# Stopping Rules for Stochastic Gradient Descent via Anytime-Valid Confidence Sequences

### v3

@@ 1.3 Related Work
Time-uniform inference and confidence sequences. Time-uniform (also called anytime-valid) inference provides uncertainty bounds that remain valid simultaneously over all times and are therefore robust to continuous monitoring and optional stopping. A foundational tool is Ville’s inequality for nonnegative supermartingales [ville1939], which converts supermartingale constructions into time-uniform deviation guarantees. Early statistical instances of this idea include classical confidence sequences for basic parameters [darling1967confidence], as well as boundary-crossing results for martingales [robbins1970boundary]. Modern developments systematize these principles through nonnegative supermartingale constructions and, in particular, mixture and stitching arguments, yielding sharp time-uniform concentration inequalities that adapt to variance-type processes [howard2021time]; see also related time-uniform Chernoff bounds based on exponential supermartingales [howard2020time] and the broader literature on self-normalized processes [de2009self]. More recently, these ideas have been reframed through e-values and game-theoretic statistics, providing a unified perspective on optional-stopping-safe inference and a comprehensive entry point to the literature [ramdas2023game, ramdas2025hypothesis]. While this framework has been developed primarily for sequential testing and estimation, it has begun to influence learning and decision-making problems, including anytime-valid inference for bandits and betting-based estimation [waudby2024anytime, waudby2024estimating]. Our work leverages this time-uniform machinery in an optimization context, using supermartingale mixtures to construct suboptimality certificates that remain valid under arbitrary, data-dependent stopping.
Stopping rules and certificates in optimization. Stopping criteria play a central practical role in stochastic optimization, yet they remain poorly understood from a statistical perspective. In practice, stochastic gradient methods are typically terminated using heuristic rules, such as monitoring validation loss, gradient norms, iterate differences, or pre-assigned computational budgets. While these rules are often effective empirically, they lack formal guarantees: they do not provide statistically valid certificates of $\varepsilon$-optimality, and their use under adaptive monitoring is generally incompatible with fixed-horizon analyses. Classical convergence results for SGD (whether asymptotic or nonasymptotic) are typically stated either in expectation or at a predetermined iteration horizon, and therefore do not justify early stopping based on observed trajectories. As a result, there is a gap between the widespread use of early stopping in optimization practice and the lack of rigorous, time-uniform performance certificates. The confidence-sequence framework developed in this paper directly addresses this gap by providing a fully data-dependent stopping rule with provable $\varepsilon$-optimality guarantees that remain valid under arbitrary, adaptive monitoring.
@@
Closest related work. The works most closely related to ours study anytime or time-uniform guarantees for stochastic approximation, but differ substantially in assumptions and in the form of the resulting guarantees. Classical results such as [rakhlin2012making] establish high-probability convergence bounds for SGD under strong convexity and bounded stochastic gradients, with guarantees stated uniformly over iterations up to a pre-specified horizon. As such, they do not support data-dependent stopping times. More recent work [feng2023anytime] derives anytime high-probability bounds for stochastic gradient methods with momentum under convexity and smoothness assumptions, using algorithm-specific dynamics and stepsize choices motivated by continuous-time ODE analysis. While these results control the optimization error uniformly over time, they do not provide explicit constants or trajectory-observable confidence sequences, and therefore cannot be used to derive certified stopping rules. Finally, [yu2025root] studies one-dimensional stochastic root finding and proposes a bisection-based procedure that yields time-uniform confidence intervals for the unknown root, but their guarantees concern inference on the unknown root rather than certificates of optimization suboptimality for projected SGD. In contrast to these works, our results apply to plain projected SGD under general convexity, yield fully explicit and trajectory-dependent confidence sequences for suboptimality, and lead to provably correct stopping rules that remain valid under arbitrary adaptive monitoring.
@@ 1.4 Organization and notation
Organization. Section 2 develops our anytime-valid confidence sequence for the weighted average suboptimality of projected SGD. Section 3 introduces the associated $\varepsilon$-optimal stopping rule and analyzes its stopping-time complexity under standard stepsize schedules. Technical lemmas and auxiliary proofs are collected in Appendix A.
Notation. For a vector $\in\mathbb{R}^{d}, x\|x\|$ denotes the Euclidean norm. For a closed convex set $\mathcal{X}\subset\mathbb{R}^{d}$, $\Pi_{\mathcal{X}}(z)$ is the Euclidean projection of $z$ onto $\mathcal{X}$. We write $\langle x,y\rangle$ for the standard inner product. Given a filtration $\{\mathcal{F}_{t}\}$, a sequence $\{X_{t}\}$ is $\mathcal{F}_{t}$-adapted if $X_{t}$ is $\mathcal{F}_{t}$-measurable, and is predictable if $X_{t}$ is $\mathcal{F}_{t-1}$-measurable. Expectations conditional on $\mathcal{F}_{t-1}$ are written $\mathbb{E}[\cdot\mid\mathcal{F}_{t-1}]$. For sequences $\{\eta_{t}\}$ and $\{x_{t}\}$, we write $S_{t}:=\sum_{s=1}^{t}\eta_{s}$ and $V_{t}:=\sum_{s=1}^{t}\eta_{s}^{2}$, and use $\bar{x}_{t}:=S_{t}^{-1}\sum_{s=1}^{t}\eta_{s}x_{s}$ to denote the weighted average iterate. Throughout, $f^{\star}:=f(x^{\star})$ denotes the optimal value and $\bar{F}_{t}$ the weighted average suboptimality defined in (4). Probabilities and expectations are taken with respect to the randomness in the stochastic gradients.
@@ 2 Anytime-Valid Confidence Sequences
where $\sigma^{2}>0$ is the sub-Gaussian variance proxy appearing in Assumption 1.1(iii). The quantity $\Sigma_{t,\mathrm{eff}}^{2}$ will be referred to as the effective cumulative variance proxy. It coincides with the true cumulative variance proxy $\Sigma_{t}^{2}$ once the latter exceeds a fixed threshold, and serves to prevent small-variance degeneracies in the time-uniform concentration argument. Larger values of $\Sigma_{t,\mathrm{eff}}^{2}$ correspond to trajectories with larger stepsizes or larger distances to the optimum, and therefore lead to wider confidence sequences. Although $Z_{t}$ is not directly observable, it is always bounded by $R_{x}^{2}$, and keeping it explicit highlights which components of the trajectory influence the final bound. A fully observable version of the confidence sequence is provided in Corollary 2.4 by replacing $Z_{t}$ with $R_{x}^{2}$.
@@
$U_{t}(\alpha):=\frac{1}{2S_{t}}\Bigg(C\,\sqrt{\Sigma_{t,\mathrm{eff}}^{2}\Big(\log\tfrac{2}{\alpha}+\log\log(\mathrm{e}+\Sigma_{t,\mathrm{eff}}^{2})\Big)}+Z_{1}+\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}\Bigg),$ (9)
where $C=C(\alpha)>0$ is an explicit constant, provided below. This boundary combines three sources of variability: the effective cumulative variance proxy $\Sigma_{t,\mathrm{eff}}^{2}$, the initial distance $Z_{1}$, and the accumulated squared gradient steps $\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}$. The leading term involving $\Sigma_{t,\mathrm{eff}}^{2}$ captures the stochastic fluctuations of the process and stems from a time-uniform concentration inequality applied to an adapted sequence satisfying conditional sub-Gaussian tail bounds (see the proof of Theorem 2.1). In contrast, the terms $Z_{1}$ and $\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}$ come directly from the deterministic structure of projected SGD, and appear when the distance recursion is telescoped. Before proceeding, we specify the constant $C$ appearing in $U_{t}(\alpha)$:
$C\;:=\;6\sqrt{\max\!\left\{1,\;\frac{\log\Big(\log\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}+1\Big)}{\log\Big(\log(\mathrm{e}+1)\Big)}\right\}},$ (10)
with $\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}$ defined as follows
$V_{\infty}\coloneqq\sum_{s=1}^{\infty}\eta_{s}^{2},\qquad\widetilde{\Sigma}_{\infty}^{2}:=\sigma^{2}R_{x}^{2}V_{\infty},\qquad\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}:=\max\!\left\{\widetilde{\Sigma}_{\infty}^{2},\;2\Big(\log\tfrac{2}{\alpha}+1\Big)\right\}.$ (11)
Since $V_{\infty}<\infty$, it follows that $\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}<\infty$, ensuring that the constant $C$ is finite. With the definition of $C$, the confidence sequence $U_{t}(\alpha)$ in (9) is now fully specified. The next theorem establishes that $U_{t}(\alpha)$ forms an anytime-valid upper confidence sequence for the weighted suboptimality process $\{\bar{F}_{t}\}_{t\geq 1}$.
@@
Suppose Assumptions 1.1 and 1.2 hold. Then, for every $\alpha\in(0,\tfrac{2}{\mathrm{e}})$, the process $\{U_{t}(\alpha)\}_{t\geq 1}$ defined in (9), with constant $C$ defined in (10), satisfies
@@
A noteworthy feature of the bound in Theorem 2.1 is its rate-controlled decay. Under the stochastic approximation stepsize conditions in Assumption 1.2, we have $S_{t}\to\infty$ and $\Sigma_{t,\mathrm{eff}}^{2}$ uniformly bounded, since the trajectory remains in a compact set and $\sigma_{t}^{2}=\sigma^{2}\eta_{t}^{2}Z_{t}$ is summable. Moreover, by Lemma A.1, the partial sums $\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}$ converge almost surely to a finite random limit. Consequently, we have the following result.
Proposition 2.2 (Rate-controlled almost-sure decay of the confidence sequence).
Under the conditions of Theorem 2.1, for any fixed $\alpha\in(0,\tfrac{2}{\mathrm{e}})$,
$U_{t}(\alpha)=O\!\left(\frac{1}{S_{t}}\right)\quad\text{almost surely},$
and in particular $U_{t}(\alpha)\to 0$ almost surely as $t\to\infty$.
This proposition ensures that projected SGD, equipped with our anytime-valid bound, provably converges to the optimum with a data-dependent, high-probability certificate of progress.
Remark 2.3 (Optimality of the $1/S_{t}$ rate).
The $1/S_{t}$ rate in Theorem 2.1 cannot, in general, be improved, even in a very simple one-dimensional deterministic setting. Fix any stepsizes $\{\eta_{t}\}_{t\geq 1}$ with $\eta_{t}>0$, $\sum_{t}\eta_{t}=\infty$, and $\sum_{t}\eta_{t}^{2}<\infty$, and consider the quadratic model
@@
with minimizer $x^{\star}=0$, $f^{\star}=0$, and deterministic oracle $g_{t}=\nabla f(x_{t})=\mu x_{t}$. Starting from $x_{1}=x_{0}\neq 0$, the projected SGD recursion reduces to the deterministic update $x_{t+1}=x_{t}-\eta_{t}g_{t}$, and the weighted cumulative suboptimality
$A_{t}:=\sum_{s=1}^{t}\eta_{s}\Big(f(x_{s})-f^{\star}\Big)$
@@
Thus, even in this benign one-dimensional model, no anytime-valid confidence sequence can shrink uniformly faster than order $1/S_{t}$. This example illustrates that the $1/S_{t}$ rate in our upper bound is essentially optimal from a time-uniform perspective. $\clubsuit$
Corollary 2.4 (Observable anytime-valid suboptimality bound).
Under the conditions of Theorem 2.1, suppose in addition that $\mathcal{X}$ is known to satisfy $\|x-x^{\star}\|\leq R_{x}$ for all $x\in\mathcal{X}$. Define
Then for every $\alpha\in(0,\tfrac{2}{\mathrm{e}})$ the process

### v4

@@ 1.3 Related work
Time-uniform inference and confidence sequences. Time-uniform (also called anytime-valid) inference provides uncertainty bounds that remain valid simultaneously over all times and are therefore robust to continuous monitoring and optional stopping. A foundational tool is Ville’s inequality for nonnegative supermartingales Ville [1939], which converts supermartingale constructions into time-uniform deviation guarantees. Early statistical instances of this idea include classical confidence sequences for basic parameters Darling and Robbins [1967], as well as boundary-crossing results for martingales Robbins and Siegmund [1970]. Modern developments systematize these principles through nonnegative supermartingale constructions and, in particular, mixture and stitching arguments, yielding sharp time-uniform concentration inequalities that adapt to variance-type processes Howard et al. [2021]; see also related time-uniform Chernoff bounds based on exponential supermartingales Howard et al. [2020] and the broader literature on self-normalized processes De la Pena et al. [2009]. More recently, these ideas have been reframed through e-values and game-theoretic statistics, providing a unified perspective on optional-stopping-safe inference and a comprehensive entry point to the literature Ramdas et al. [2023]; Ramdas and Wang [2025]. While this framework has been developed primarily for sequential testing and estimation, it has begun to influence learning and decision-making problems, including anytime-valid inference for bandits and betting-based estimation Waudby-Smith et al. [2024]; Waudby-Smith and Ramdas [2024]. Our work leverages this time-uniform machinery in an optimization context, using supermartingale mixtures to construct performance certificates for stochastic gradient methods that remain valid uniformly over time under arbitrary, data-dependent stopping.
Stochastic approximation and asymptotic normality. Stochastic approximation provides the theoretical foundation for stochastic gradient methods. The seminal paper Robbins and Monro [1951] established almost sure convergence of iterative schemes under diminishing stepsizes, while Kushner and Yin [2003] developed a comprehensive asymptotic theory based on ordinary differential equations, covering stability, asymptotic normality, and efficiency. A key advance showed that appropriately averaged iterates achieve asymptotically optimal covariance, providing a principled justification for studying weighted averages (which we adopt in this paper) of SGD trajectories Polyak and Juditsky [1992]. This asymptotic perspective was sharpened in Duchi and Ruan [2021], which characterized minimax and asymptotically efficient rates for stochastic optimization, including constrained formulations, which was further extended to nonsmooth stochastic approximation in Davis et al. [2024]. Together, these results motivate the use of averaged iterates and enable asymptotic confidence regions for optimal solutions based on limiting normality. However, such guarantees are inherently asymptotic: they do not yield finite-time confidence sets, nor do they provide time-uniform validity under continuous monitoring or adaptive stopping.
Fixed-horizon optimality guarantees for convex SGD. A vast literature develops nonasymptotic convergence guarantees for stochastic gradient (and mirror-descent–type) methods that hold with high probability at a pre-specified iteration horizon. Early robust frameworks for stochastic convex optimization established high-probability bounds under bounded or sub-Gaussian noise assumptions, providing optimal oracle complexity guarantees for stochastic approximation and stochastic programming Nemirovski et al. [2009]; Lan [2012]. Related work also studied statistical validation of mirror-descent–type stochastic approximation methods, providing confidence bounds for the optimal objective value at a fixed iteration horizon rather than under adaptive or continuously monitored stopping rules Lan et al. [2012]. Parallel developments in machine learning emphasized refined nonasymptotic analyses of stochastic approximation algorithms, including explicit bias–variance tradeoffs and finite-time rates for averaged iterates Moulines and Bach [2011]; Bach and Moulines [2013]; Dieuleveut and Bach [2016]. Subsequent work sharpened fixed-horizon high-probability bounds across a range of regimes, including nonsmooth and strongly convex objectives, and clarified when optimal rates can be achieved by averaging versus last-iterate solutions Rakhlin et al. [2012]; Shamir and Zhang [2013]; Harvey et al. [2019a, b]; Jain et al. [2021]. Complementary information-theoretic lower bounds characterize the fundamental limits of stochastic convex optimization at fixed confidence levels Agarwal et al. [2012]. More recent contributions extend fixed-horizon high-probability guarantees beyond light-tailed noise models, including heavy-tailed settings handled via gradient clipping and related robustification techniques Gorbunov et al. [2020], as well as general boosting and robust-estimation wrappers that convert low-probability guarantees into high-confidence ones Davis et al. [2021]. Together, these works provide sharp finite-time guarantees at prescribed horizons and confidence levels. However, these guarantees are intrinsically fixed-time: they are stated for pre-specified horizons and do not remain valid under continuous monitoring or data-dependent stopping without additional union bounds or restarts.
Fixed-horizon stationarity guarantees for nonconvex SGD. Theoretical guarantees for stochastic gradient methods in nonconvex optimization have been studied extensively, with performance typically quantified through first-order stationarity criteria, most commonly bounds on the squared gradient norm $\|\nabla f(x)\|^{2}$ evaluated at a prescribed iteration horizon via a suitable output rule. Foundational stochastic approximation theory establishes almost sure convergence of iterates to stationary sets under diminishing stepsizes and appropriate noise conditions, providing the conceptual basis for modern analyses in nonconvex stochastic optimization Robbins and Monro [1951]; Kushner and Yin [2003]; Benaïm [2006]; Borkar and Borkar [2008]. For smooth nonconvex objectives, a widely adopted nonasymptotic benchmark is an $O(t^{-1/2})$ rate for approximate stationarity under unbiased stochastic gradients and bounded-variance or light-tailed noise. Such guarantees are typically stated in expectation for a prescribed output iterate, such as a randomly selected or averaged iterate among the first $t$ iterates, or equivalently in terms of the minimum stationarity measure attained up to time $t$. This benchmark was crystallized and systematically analyzed in influential works Ghadimi and Lan [2013, 2016], and has since served as a reference point for a broad range of subsequent contributions; see also modern syntheses such as Bottou et al. [2018]; Lan [2020]. Beyond vanilla SGD, extensive work studies refined oracle-complexity guarantees for nonconvex problems, including variance-reduced and recursive-gradient methods for finite-sum or stochastic objectives Reddi et al. [2016]; Lei and Jordan [2017]; Nguyen et al. [2017]; Fang et al. [2018]; Li et al. [2021], algorithms designed to escape saddle points Ge et al. [2015]; Jin et al. [2017]; Allen-Zhu [2018], and adaptive or momentum-based schemes in nonconvex landscapes Zaheer et al. [2018]; Ward et al. [2020]. Complementary results establish nonasymptotic stationarity guarantees in nonconvex optimization through different mechanisms, including high-probability bounds obtained via gradient clipping or other robustification techniques under heavy-tailed noise Cutkosky and Mehta [2021]; Li and Liu [2022], adaptive first-order methods with refined concentration analyses Kavis et al. [2022], and expectation-based analyses of vanilla SGD under heavy-tailed stochastic gradients Fatkhullin et al. [2025]. While this body of work provides sharp convergence and oracle-complexity guarantees at fixed, pre-specified horizons, it does not address the statistical validity of adaptive stopping decisions based on continuously monitored optimization trajectories.
@@
Closest related work. The works most closely related to ours study anytime or time-uniform guarantees for stochastic approximation, but differ substantially in assumptions and in the form of the resulting guarantees. Classical results such as Rakhlin et al. [2012] establish high-probability convergence bounds for SGD under strong convexity and bounded stochastic gradients, with guarantees stated uniformly over iterations up to a pre-specified horizon. As such, they do not support data-dependent stopping times. More recent work Feng et al. [2023] derives anytime high-probability bounds for stochastic gradient methods with momentum under convexity and smoothness assumptions, using algorithm-specific dynamics and stepsize choices motivated by continuous-time ODE analysis. However, these results are not formulated in terms of trajectory-observable confidence sequences and are not used to derive certified stopping rules. Finally, Yu et al. [2025] studies one-dimensional stochastic root finding and proposes a bisection-based procedure that yields time-uniform confidence intervals for the unknown root, but their guarantees concern inference on the unknown root rather than certificates of optimization suboptimality for projected SGD. In contrast to these works, our results apply to plain projected SGD under general convexity, yield fully explicit and trajectory-dependent confidence sequences for suboptimality, and lead to provably correct stopping rules that remain valid under arbitrary adaptive monitoring.
@@ 1.4 Organization and notation
Organization. Section 2 develops anytime-valid upper confidence sequences for the weighted average suboptimality in the convex case and for a weighted stationarity measure in the nonconvex case, including fully observable variants, and studies their decay rates under common stepsize schedules. Section 3 uses these confidence sequences to construct statistically valid, trajectory-dependent stopping rules and analyzes the resulting stopping-time complexity under common stepsize schedules. Section 4 contains the proofs of the main results. Finally, the Appendix collects additional proofs and supporting technical lemmas.
Notation. For a vector $\in\mathbb{R}^{d}, x\|x\|$ denotes the Euclidean norm and $\langle x,y\rangle$ the standard inner product. For a closed convex set $\mathcal{X}\subset\mathbb{R}^{d}$, $\Pi_{\mathcal{X}}(z)$ denotes the Euclidean projection of $z$ onto $\mathcal{X}$. Given a filtration $\mathcal{F}_{t}$}, a sequence $\{X_{t}\}$ is $\{\mathcal{F}_{t}\}$-adapted if $X_{t}$ is $\mathcal{F}_{t}$-measurable, and is predictable if $X_{t}$ is $\mathcal{F}_{t-1}$-measurable. Conditional expectations are written $\mathbb{E}[\cdot\mid\mathcal{F}_{t-1}]$. We write $\inf\varnothing=\infty$. For two nonnegative sequences $\{a_{t}\}$ and $\{b_{t}\}$, we write $a_{t}=O(b_{t})$ if there exists a constant $C<\infty$ such that $a_{t}\leq C,b_{t}$ for all $t$; we write $a_{t}\lesssim b_{t}$ synonymously with $a_{t}=O(b_{t})$, and $a_{t}\asymp b_{t}$ if $a_{t}\lesssim b_{t}$ and $b_{t}\lesssim a_{t}$. Unless stated otherwise, implicit constants in $O(\cdot)$, $\lesssim$, and $\asymp$ may depend on fixed problem parameters but are independent of the time index $t$ (and independent of $\varepsilon$ in stopping-time bounds). Throughout, probabilities and expectations are taken with respect to the randomness in the stochastic gradients.
@@ 2 Anytime-Valid Confidence Sequences
In this section we develop anytime-valid confidence sequences for stochastic gradient descent that remain valid under continuous monitoring and arbitrary data-dependent stopping. The constructions are trajectory-adaptive, depending on quantities generated along the SGD path rather than on fixed global problem parameters. We treat the convex and nonconvex settings separately. In the convex case, we derive confidence sequences for weighted average suboptimality of projected SGD. In the nonconvex case, we obtain time-uniform certificates for weighted first-order stationarity. Both constructions admit fully observable variants suitable for online use.
2.1 Convex case: suboptimality certificates
@@
where $\sigma^{2}>0$ is the sub-Gaussian variance proxy from Assumption 1.2(ii). The quantity $\Sigma_{t,\mathrm{eff}}^{2}$ will be referred to as the effective cumulative variance proxy. It coincides with the true cumulative variance proxy $\Sigma_{t}^{2}$ once the latter exceeds a fixed threshold, and serves to prevent small-variance degeneracies in the time-uniform concentration argument. Larger values of $\Sigma_{t,\mathrm{eff}}^{2}$ correspond to trajectories with larger stepsizes or larger distances to the optimum, and therefore lead to wider confidence sequences. Although $Z_{t}$ is not directly observable, it is always bounded by $R_{x}^{2}$, and keeping it explicit highlights which components of the trajectory influence the final bound. A fully observable version of the confidence sequence is provided in Corollary 2.2 by replacing $Z_{t}$ with $R_{x}^{2}$.
@@
$U_{t}(\alpha):=\frac{1}{2S_{t}}\Bigg(7\,\sqrt{\Sigma_{t,\mathrm{eff}}^{2}\Big(\log\tfrac{2}{\alpha}+\log\log(\mathrm{e}+\Sigma_{t,\mathrm{eff}}^{2})\Big)}+Z_{1}+\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}\Bigg).$ (6)
This boundary combines three sources of variability: the effective cumulative variance proxy $\Sigma_{t,\mathrm{eff}}^{2}$, the initial distance $Z_{1}$, and the accumulated squared gradient steps $\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}$. The leading term involving $\Sigma_{t,\mathrm{eff}}^{2}$ captures the stochastic fluctuations of the process and stems from a time-uniform concentration inequality applied to an adapted sequence satisfying conditional sub-Gaussian tail bounds (see the proof of Theorem 2.1). In contrast, the terms $Z_{1}$ and $\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}$ come directly from the structure of projected SGD, and appear when the distance recursion is telescoped. The next theorem establishes that $U_{t}(\alpha)$ forms an anytime-valid upper confidence sequence for the weighted suboptimality process $\{\bar{F}_{t}\}_{t\geq 1}$.
@@
Suppose Assumptions 1.1 and 1.2 hold. Then, for every $\alpha\in(0,\tfrac{2}{\mathrm{e}})$, the process $\{U_{t}(\alpha)\}_{t\geq 1}$ defined in (6) satisfies
@@
Corollary 2.2 (Observable anytime-valid suboptimality bound).
Under the conditions of Theorem 2.1, suppose in addition that $\mathcal{X}$ satisfies $\|x-x^{\star}\|\leq R_{x}$ for all $x\in\mathcal{X}$. Define
Then for every $\alpha\in(0,\tfrac{2}{\mathrm{e}})$ the process
$U_{t}^{\mathrm{obs}}(\alpha):=\frac{1}{2S_{t}}\Bigg(7\sqrt{\widetilde{\Sigma}_{t,\mathrm{eff}}^{2}\!\left(\log\tfrac{2}{\alpha}+\log\log(\mathrm{e}+\widetilde{\Sigma}_{t,\mathrm{eff}}^{2})\right)}+R_{x}^{2}+\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}\Bigg)$
Note that $U_{t}^{\mathrm{obs}}(\alpha)$ is $\mathcal{F}_{t}$-measurable and can be computed online at each iteration using only observable information. The following remark shows that the $1/S_{t}$ normalization appearing in the bound is unavoidable from a time-uniform perspective, even in simple deterministic settings. While the numerator in $U_{t}(\alpha)$ may exhibit different behavior depending on the stepsize regime and the realized stochastic gradients, no anytime-valid confidence sequence can shrink uniformly faster than order $1/S_{t}$ in general.
Remark 2.3 (Unavoidability of the $1/S_{t}$ normalization).
The $1/S_{t}$ normalization in Theorem 2.1 is, in general, unavoidable from a time-uniform (anytime-valid) perspective, even in a very simple one-dimensional deterministic setting. Fix any stepsizes $\{\eta_{t}\}_{t\geq 1}$ with $\eta_{t}>0$, $\sum_{t}\eta_{t}=\infty$, and $\sum_{t}\eta_{t}^{2}<\infty$, and consider the quadratic model
@@
with minimizer $x^{\star}=0$, $f(x^{\star})=0$, and deterministic oracle $g_{t}=\nabla f(x_{t})=\mu x_{t}$. Starting from $x_{1}=x_{0}\neq 0$, the projected SGD recursion reduces to the deterministic update $x_{t+1}=x_{t}-\eta_{t}g_{t}$, and the weighted cumulative suboptimality
$A_{t}:=\sum_{s=1}^{t}\eta_{s}\Big(f(x_{s})-f(x^{\star})\Big)$
@@
Thus, even in this benign one-dimensional model, time-uniform validity forces $U_{t}(\alpha)$ to scale no better than order $1/S_{t}$ along this path. This example illustrates that the appearance of the factor $1/S_{t}$ in time-uniform (anytime-valid) upper bounds is essentially necessary. $\clubsuit$
The following remark clarifies that the assumption $\|x-x^{\star}\|\leq R_{x}$ for all $x\in\mathcal{X}$ is sufficient but not necessary for constructing an observable confidence sequence.
Remark 2.4 (Compactness of $\mathcal{X}$).
The compactness of $\mathcal{X}$ is not required for the validity of Theorem 2.1, which holds without any boundedness assumption on the iterates. Compactness is used only to obtain the observable confidence sequence in Corollary 2.2, where a uniform bound $R_{x}$ on the distances $\|x_{t}-x^{\star}\|$ along the SGD trajectory is needed. Projection onto a compact set provides a simple and explicit mechanism for ensuring such a bound. However, both compactness and projection can be omitted whenever an a priori bound $\|x_{t}-x^{\star}\|\leq R_{x}$ is known to hold along the trajectory, for instance under stability conditions that guarantee the iterates remain in a bounded region. $\clubsuit$
2.2 Nonconvex case: stationarity certificates
In this subsection we extend the confidence-sequence framework to nonconvex SGD, where progress is quantified through first-order stationarity. Throughout this section, we consider unconstrained optimization over $\mathcal{X}=\mathbb{R}^{d}$ with a differentiable, globally smooth, and nonnegative objective, so that the SGD recursion takes the standard form without projection. For a given confidence level $\alpha\in(0,1)$, our goal is to construct a time-uniform upper confidence sequence $\{W_{t}(\alpha)\}_{t\geq 1}$ for the weighted average stationarity measure $\bar{G}_{t}$ defined in (4). By Definition 1.4, this ensures that the bound $\bar{G}_{t}\leq W_{t}(\alpha)$ holds simultaneously for all $t\geq 1$ with probability at least $1-\alpha$, even under continuous monitoring and adaptive stopping.
As in the convex case, the construction is trajectory-adaptive, but the source of adaptivity differs. In the nonconvex setting, the width of the confidence sequence is driven by the realized gradient norms rather than distances to an optimizer. This reflects the geometry of nonconvex optimization, where stationarity replaces optimality as the relevant notion of progress. Consequently, the certificate tightens automatically when SGD enters low-gradient regions and becomes more conservative along trajectories exhibiting larger stochastic variation or larger gradients.
To state the main result, we introduce a trajectory-dependent variance proxy that governs the width of the confidence sequence. For each $t\geq 1$, define